OpenStudy (anonymous):

how to solve 3 x 3^root(8)=x

OpenStudy (anonymous):

\[3 \times \sqrt[3]{8} = 8\]

OpenStudy (misty1212):

lol no it isn't silly boy

OpenStudy (anonymous):

We can try to approach this question, by first understanding the powers of 2, to get 8

OpenStudy (misty1212):

\[\sqrt[3]{8}=2\] so \[3\times 2=6\] not \(8\) doe

OpenStudy (anonymous):

\[2^{3}= 8\] you mean that?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so the asnwer is 8/3 ?

OpenStudy (anonymous):

No, you can't just divide 8 by 3

OpenStudy (anonymous):

When you move powers to the other side, you do the opposite of powers, which is nth rooting the value

OpenStudy (misty1212):

lol you just making this stuff up right?

OpenStudy (anonymous):

I accidentally typed it in the wrong comment section ><

OpenStudy (jack1):

lol